OpenStudy (anonymous):

(2√5+3√7)^2 Can someone please show me and explain what your doing to get it??

OpenStudy (anonymous):

here you have to use this formula: \[\LARGE (a+b)^2=a^2+2ab+b^2\] so \[\LARGE (2\sqrt5+3\sqrt7)^2=?\] can you apply the formula? :)

OpenStudy (anonymous):

@mojoman5465

OpenStudy (anonymous):

I've tried doing it but Im really confused on how to do it thats whey i asked..

OpenStudy (anonymous):

\[\LARGE (2\sqrt5)^2+2[\cdot2\sqrt5\cdot 3\sqrt7]+(3\sqrt7)^2=\] can you continue any further? :)

OpenStudy (anonymous):

\[\LARGE 4\cdot 5+12\sqrt{5\cdot 7}+9\cdot 7 =...\]